OpenStudy (anonymous):

solve for b: (x/a)+(y/b)+(z/c)=1

OpenStudy (anonymous):

\[1-\frac{z}{c}-\frac{x}{a}=\frac{y}{b}\] Now get the inverse of both sides: \[\frac{b}{y}=\frac{1}{1-\frac{z}{c}-\frac{x}{a}}\] Times both sides by y: \[b=\frac{y}{1-\frac{z}{c}-\frac{x}{a}}\]